5x+y/1.3y+6=
x=.24 y=.5
please solve

Answer by Eric Johnston(7)

My first observation is that the two y's cancel out so the input value of y is irrelevant; so maybe the question is supposed to be
5x+y/{1.3y+6}=
x=.24 y=.5
Substitute values of x=0.24 and y=0.5 into equation:
Ans = 5 * 0.24 + 0.5/{1.3 * 0.5 + 6}
Ans = 1.2 + 0.5/{0.65 + 6}
Ans = 1.2 + 0.5/6.65
Ans = 1.2 + 0.075
Ans = 1.275
I hope that is right. It is the first time I have answered one of these questions.